# Effective synchronization amid noise-induced chaos

### v1

@@ III Convergent distributions
To this end, we simulate two oscillator ensembles, each consisting of $N=500$ phase samples. Their phase values are randomly and independently selected from different uniform distributions; $\{\varphi^{n}_{\mathrm{A}}\sim\mathrm{U}[0,u)\}$ for the sender and $\{\varphi^{n}_{\mathrm{B}}\sim\mathrm{U}[0.5,0.5+u)\}$ for the receiver. By this, we study the extreme case of well-separated (non-overlapping) initial phase distributions. We subject the oscillators to Eq. (1) for given Lyapunov exponents $\Lambda$ and initial widths $u$. All $2N$ phases evolve under the same $\psi(\varphi)$ and uniformly-distributed waiting times $\{\beta^{(k)}\sim\mathrm{U}[0,1)\}$. At each kick, we estimate $S^{(k)}_{\mathrm{A}}$, $S^{(k)}_{\mathrm{B}}$, $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$, and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$. [^4] The KLD is not symmetric. However, since the initial distributions are offset by exactly half of the phase circle and $\beta\in[0,1)$, we expect both $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ to have the same statistics over many repetitions. Indeed their averages of many realizations, depicted in Fig. 4(a,b), are equal.
For completeness, in SI Appendix, Sec. S5 we show how the synchronization for $\Lambda<0$ manifests in the entropy (cf. also Ref. [21]) and the KLD. The KLD has not been considered previously in the literature, as only a single ensemble has been considered until present.
@@
For $\Lambda>0$, the two distributions $q_{\mathrm{A}}^{(k)}(\varphi)$ and $q_{\mathrm{B}}^{(k)}(\varphi)$ change erratically between consecutive kicks, e.g., as peaks are formed at phase values in the vicinity of the phase map’s extrema, and ‘smeared’ away from them (see Fig. 2 and SI Appendix, Fig. S4). Accordingly, in Fig. 3(b), we see that the entropy also changes chaotically [21]. At the same time, remarkably, both KLDs decay to zero as seen in Fig. 3(a), indicating that the two ensembles eventually converge to the same distribution. [^5] The brief transient during $25\lesssim k\lesssim 65$, where the KLD seems to stay at a constant value is a result of the distributions being very narrow at these time steps (see Fig. 2), so they change widths together without affecting the KLD. We quantitatively explain this behavior in SI Appendix, Sec. S5. When the distribution widen afterwards, the KLD decreases to below its numerical error cutoff. Accordingly, the entropies of both distributions coincide at later times in Fig. 3(b). Note that a different realization of waiting times $\{\beta^{(k)}\}$ would produce vastly different distributions (see SI Appendix, Fig. S5); what we report, therefore, is that the two agents’ distributions agree more closely with more kicks for every stochastic realization.
We qualitatively understand this ensemble convergence as arising from independence of initial state. After many kicks, Eq. (1) approaches a chaotic map from the phase circle onto itself with Lyapunov exponent $\Lambda$ per kick. Consider the map from an initial phase $\varphi^{(0)}$ to the final phase $\varphi^{(k)}$ for a particular sequence of random and uniformly-distributed waiting times $\{\beta^{(k)}\}$. This map comes from a statistically uniform sampling of the phase map $\psi(\varphi)$ and has an overall Lyapunov exponent for any $\varphi^{(0)}$ that approaches $k\Lambda$. Any such map fully mixes all $\varphi^{(0)}$[33]. That is, the $\varphi^{(k)}$ for a given $\varphi^{(0)}$ may be arbitrarily close to the $\varphi^{\prime(k)}$ for other $\varphi^{\prime(0)}$’s residing anywhere throughout the phase circle. Thus, any uncertainty in an observed $\varphi^{(k)}$ erases all knowledge of $\varphi^{(0)}$. So, we expect that for a given uncertainty in $\varphi^{(k)}$ there exists some kick number $K_{\mathrm{m}}$ such that any dependence of $\varphi^{(k)}$ on $\varphi^{(0)}$ for $k>K_{\mathrm{m}}$ is undetectable. This diminishing uncertainty manifests, therefore, in a KLD that approaches zero, implying that the two distributions differ only by this decaying uncertainty. We proceed to study the rate with which the oscillators become statistically indistinguishable.
@@
Figure 4: The average Kullback-Leibler divergence (KLDs) $D({\mathrm{A}}\|{\mathrm{B}})$ and $D({\mathrm{B}}\|{\mathrm{A}})$ among the two distributions, averaged over $500$ waiting-time realizations. We used phase maps of Eq. (2) with various values of $\Lambda$ obtained by varying the gain parameter $A$. We show the results for (a) $\Lambda=0.405$ and (b) $\Lambda=0.121$. The number of phase samples is $N=500$ in each ensemble, and they are initially uniform-distributed with width $u=10^{-4}$. Dashed orange lines depict the exponential fit from which we extract the mixing-kick number. (c) The mixing-kick number $K_{\mathrm{m}}$ versus Lyapunov exponent $\Lambda$. $K_{\mathrm{m}}$ is identified as minus the inverse slope in the exponential decay of $\langle D\rangle$. The exponential regime begins after a transient where the distributions spread and overlap, and terminates when the KLD is comparable with its numerical estimation error, $\sim 0.05$. Mixing occurs faster as $\Lambda$ increases, so the two agents typically converge earlier. The decay is the powerlaw $K_{\mathrm{m}}=2.82\Lambda^{-1.49}$.
@@ III.1 Convergence rate
For all $\Lambda$’s, we observe a predominant regime of an exponential decay with $k$, shown with a dashed orange lines in Fig. 4(a,b). [^6] The true KLDs should have started from $\infty$ (as the two initial distributions are nonoverlapping), but instead they begin at a large finite value, $D^{(0)}\simeq 15.1\sim{-\log u}$. This is an artifact of the numerical KLD estimate [32] (Eq. (10)), and thus the KLD values at the first few kicks in Fig. 4 should be taken with caution. The exponential decay due to mixing begins past that transient. The decay is to be expected due to the stochastic evolution of the distributions, which are mixed until becoming equivalent. This exponential decay is typically cleaner for higher $\Lambda$’s (compare $k\in(10,45)$ in Fig. 4(a) with $k\in(70,250)$ in Fig. 4(b)), and lasts until we reach the numerical error cutoff for the KLD estimator, $\sim 0.05$.
For each $\Lambda$, we fit an exponential (as in Figs. 4(a,b)) to obtain a decay rate defined as $1/K_{m}(\Lambda)$. We interpret these $K_{\mathrm{m}}$’s as the mixing-kick number. We verify for various $\Lambda$’s that it does not depend on the initial condition (by simulating $u=5\cdot 10^{-4}$ as well) or the sample size (by simulating $N=100$, too); $K_{\mathrm{m}}(\Lambda)$ remained identical within $3\%$ (not shown) for each $\Lambda$. This supports our claim that the decay is predominantly exponential, and suggests that our choice for the mixing-kick number is meaningful as it is only a property of $\Lambda$ and is independent of the numerical limitations for the class of phase maps studied.
In Fig. 4(c) we plot the mixing-kick number $K_{\mathrm{m}}$ versus $\Lambda$. As expected, the more violent the phase map (bigger $\Lambda$) the faster the mixing. We find numerically that the dependence of $K_{\mathrm{m}}$ on $\Lambda$ is consistent with a power law, $K_{\mathrm{m}}(\Lambda)\sim\Lambda^{-1.491\pm 0.046}$. We caution that the exponential fit for the decay of the KLD with low $\Lambda$ is not perfect (specifically, for $\Lambda\lesssim 0.2$), so their $K_{\mathrm{m}}$ should be taken with a grain of salt. The exact functional form of $K_{\mathrm{m}}(\Lambda)$ is not central to what follows; we shall only use the fact that $K_{\mathrm{m}}$ decays with $\Lambda$ as a qualitative guide. The power-law does give the expected limiting behavior, viz., $K_{\mathrm{m}}{(\Lambda\to 0^{+})}\to+\infty$.
To summarize, for $\Lambda>0$, the two phase statistics become identical after iterating sufficiently longer than the mixing-kick number $K_{\mathrm{m}}$, which is shown to decay as a power law with $\Lambda$. Thus, for the purpose of establishing synchronization, we will first make the reasonable assumption that both agents have received a number of noisy kicks exceeding an agreed threshold. In the next section we demonstrate this synchronization.
@@ IV Effective synchronization
In Sec. III, we found that agents ${\mathrm{A}}$ and ${\mathrm{B}}$ must eventually share $q^{(k)}(\varphi)$ in both the $\Lambda<0$ (synchronizing) and the $\Lambda>0$ (unsynchronizing) regimes. With $\Lambda<0$, the two agents agree at every moment (sufficiently-long after the most recent kick) on a specific phase to experimental accuracy [15] and can use it to perform simultaneous actions or communicate. For $\Lambda>0$, while the resultant distributions do not asymptotically approach a sharp peak, the distributions of two agents become fully identical as captured by the approach of the KLD among them to zero. The synchrony in the former and convergence in the latter remain with arbitrarily more kicks. Thus, any functional of $q(\varphi)$ calculated independently by both agents would be shared after any large $k$. Below we consider a functional we refer to as the “fiducial phase”, $\varphi_{\mathrm{f}}[q(\varphi)]$, that indicates a virtual phase position upon which both agents must agree at a given moment.
Given the value of $\varphi_{{\mathrm{f}}}^{(k)}=\varphi_{\mathrm{f}}[q^{(k)}(\varphi)]$ after a kick, any agent can infer its value at a later time $\varphi_{{\mathrm{f}}}^{(k)}+t/T$ from its own phase at that moment, as it has also advanced from $\varphi_{n}^{(k)}$ to $\varphi_{n}^{(k)}+t/T$. Since the kicks occur at common timing, the time $t$ that has passed after the kick is the same for all agents, and hence any difference between their current $\varphi_{\mathrm{f}}^{(k)}+t/T$’s is the same as the difference between their initial $\varphi_{\mathrm{f}}^{(k)}$’s, $[\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}+t/T]-[\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}+t/T]=\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}-\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$. Thus, if the prescription for $\varphi_{\mathrm{f}}^{(k)}$ is such that $\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}\simeq\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$, their $\varphi_{\mathrm{f}}^{(k)}+t/T$’s are also about equal. The two $\varphi_{\mathrm{f}}^{(k)}+t/T$ are thus effectively synchronized over time; see illustration in SI Appendix, Fig. S1. We demonstrate below that this type of synchronization can indeed be achieved.
@@
The uncertainty in $\varphi_{\mathrm{f}}$ must depend on the degree of nonuniformity of $q(\varphi)$; surely, no $\varphi_{\mathrm{f}}$ can be unambiguously determined if $q(\varphi)$ is completely uniform. Further, uncertainty in measuring $q(\varphi)$ leads to further uncertainty in $\varphi_{\mathrm{f}}$. Below we test the discrepancy between $\varphi_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi_{{\mathrm{f}},{\mathrm{B}}}$ using the finite sampling method employed above. Our proof-of-concept example below shows that with a fixed sample size $N$, it is indeed feasible to attain arbitrarily-small discrepancies which further decrease with bigger $N$. These discrepancies depend strongly on the nonuniformity of the $q(\varphi)$ at hand, as measured by its entropy, Eq. (6). Thus, reliably small discrepancies require phase maps $\psi(\varphi)$ with small typical entropies.
@@
To determine a peak position, one needs a smooth estimate of $q(\phi)$ from the sampled values $\{\varphi_{1},\ldots,\varphi_{N}\}$. We obtain an estimate denoted $Q(\varphi,\sigma)$ smoothed to a “bandwidth” $\sigma$ using the method of kernel-density estimation [34, 35, 36], detailed in Sec. A.2. The reliability of the $\varphi_{\mathrm{f}}$ obtained depends on $\sigma$. Large $\sigma$ produces a $Q(\varphi;\sigma)$ with a few broad peaks whose positions have high uncertainty, whereas small $\sigma$ gives numerous statistically insignificant peaks; see illustration in SI Appendix, Figs. S7.777Finding an optimal bandwidth $\sigma$, particularly for highly-multimodal distributions, is an open question in the field [37, 38, 39, 40].
@@
This crude choice turns out to be sufficient for our purpose.

### v2

@@ III Convergent distributions
To this end, we simulate two oscillator ensembles, each consisting of $N=500$ phase samples. Their phase values are randomly and independently selected from different uniform distributions $\varphi^{(0)}_{{\mathrm{A}}n}\in[0,u)$ for the sender and $\varphi^{(0)}_{{\mathrm{B}}n}\in[0.5,0.5+u)$ for the receiver. By this, we study the extreme case of well-separated (non-overlapping) initial phase distributions. [^4] The true KLDs should have started from $\infty$ as the two initial distributions are nonoverlapping. Instead, in Figs. 3(a) and 4(a,b) they begin at a large finite value, $D^{(0)}\sim{-\log u}$. This is an artifact of the numerical KLD estimate (Eq. (10)). The exponential decay due to the mixing we report commences past that short transient, at which point the two distributions have the same support $\varphi\in[0,1)$, wherein indeed the KLD is a reliable measure. Other measures of similarity betweeen two distributions exist — notably the Wasserstein measure [36]. It is not singular for cases of non-overlapping supports, although it is only indirectly affected by the distributions’ peak heights; a central aspect to our analysis. Our choice of the KLD measure is motivated by our information-theoretic approach in terms of entropy. We subject the oscillators to Eq. (1) for given Lyapunov exponents $\Lambda$ and initial widths $u$. All $2N$ phases evolve under the same $\psi(\varphi)$ and uniformly-distributed waiting times $\beta^{(k)}\in[0,1)$. At each kick, we estimate $S^{(k)}_{\mathrm{A}}$, $S^{(k)}_{\mathrm{B}}$, $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$, and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$. [^5] The KLD is not symmetric. However, since the initial distributions are offset by exactly half of the phase circle and $\beta\in[0,1)$, we expect both $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ to have the same statistics over many repetitions. Indeed their averages over realizations are equal in Fig. 4(a,b).
@@
For completeness, in SI Appendix, Sec. S5 we show how the synchronization for $\Lambda<0$ manifests in the entropy (see also Ref. [24]) and the KLD. For $\Lambda>0$, the two distributions $q_{\mathrm{A}}^{(k)}(\varphi)$ and $q_{\mathrm{B}}^{(k)}(\varphi)$ change erratically between consecutive kicks, e.g., as peaks are formed at phase values in the vicinity of the phase map’s extrema, and ‘smeared’ away from them (see Fig. 2 and SI Appendix, Fig. S4). Accordingly, in Fig. 3(b), we see that the entropy also changes chaotically [24]. At the same time, remarkably, both numerical KLDs decay to zero as seen in Fig. 3(a) despite the nonoverlapping supports of the initial distributions, indicating that the two ensembles eventually converge to the same distribution. [^6] The brief transient during $25\lesssim k\lesssim 65$, where the KLD seems to stay at a constant value is a result of the distributions being very narrow at these time steps (see Fig. 2), so they change widths together without affecting the KLD. We quantitatively explain this behavior in SI Appendix, Sec. S5. When the distribution widen afterwards, the KLD decreases to below its numerical error cutoff. Accordingly, the entropies of both distributions coincide at later times in Fig. 3(b). Note that a different realization of waiting times $\{\beta^{(k)}\}$ produces vastly different distributions (see SI Appendix, Fig. S5); what we report, therefore, is that the two agents’ distributions agree more closely with more kicks for every stochastic realization.
@@
Figure 4: The average Kullback-Leibler divergence (KLDs) $D({\mathrm{A}}\|{\mathrm{B}})$ and $D({\mathrm{B}}\|{\mathrm{A}})$ among the two distributions, averaged over $500$ waiting-time realizations. We used phase maps of Eq. (2) with various values of $\Lambda$ obtained by varying the gain parameter $A$. We show the results for (a) $\Lambda=0.405$ and (b) $\Lambda=0.121$. The number of phase samples is $N=500$ in each ensemble, and they are initially uniform-distributed with width $u=10^{-4}$. Dashed orange lines depict the exponential fit from which we extract the mixing-kick number. (c) The mixing-kick number $K_{\mathrm{m}}$ versus Lyapunov exponent $\Lambda$. The two enlarged orange points correspond to $K_{\mathrm{m}}$ for the $\Lambda$’s shown in panels (a) and (b). $K_{\mathrm{m}}$ is identified as minus the inverse slope in the exponential decay of $\langle D\rangle$. The exponential regime begins after a transient where the distributions spread and overlap, and terminates when the KLD is comparable with its numerical estimation error, $\sim 0.05$. Mixing occurs faster as $\Lambda$ increases, so the two agents typically converge earlier. The decay is consistent with the powerlaw $K_{\mathrm{m}}=2.82\Lambda^{-1.49}$.
We qualitatively understand this ensemble convergence as arising from independence of initial state. Consider the map from an initial phase $\varphi^{(0)}$ to the final phase $\varphi^{(k)}$ for a particular sequence of random and uniformly distributed waiting times $\{\beta^{(k)}\}$. This map comes from a statistically uniform sampling of the phase map $\psi(\varphi)$ and has an overall Lyapunov exponent for any $\varphi^{(0)}$ that approaches $k\Lambda$ as $k$ gets large. Since the phase circle is periodic and finite, nearby phases cannot be drifting apart as $\sim e^{k\Lambda}$ indefinitely. Instead, the non-monotonic phase map (Eq. (2)) “stretches and folds” distant phases akin to the Baker’s map, such that full mixing of $\varphi^{(0)}$s is possible [37]. [^7] This qualitative argument may suggest that full mixing is inevitable for such maps, though we know of neither a rigorous proof nor the necessary conditions for it to hold. Here, we have merely verified that the convergence in our examples was sufficient to produce closely-agreeing fiducial phases. As a result, the $\varphi^{(k)}$ for a given $\varphi^{(0)}$ may be arbitrarily close to the $\varphi^{\prime(k)}$ for other $\varphi^{\prime(0)}$’s residing anywhere throughout the phase circle. Then, any uncertainty in an observed $\varphi^{(k)}$ would erase all knowledge of $\varphi^{(0)}$. So, we expect that for any given uncertainty in $\varphi^{(k)}$ there exists some kick number $K_{\mathrm{m}}$ such that any dependence of $\varphi^{(k)}$ on $\varphi^{(0)}$ for $k>K_{\mathrm{m}}$ is undetectable [38]. This diminishing detectability would entail a KLD that approaches zero, implying that the two distributions differ only by this decaying uncertainty. We proceed to systematically show that the KLD in our examples indeed converges to zero.
@@ III.1 Convergence rate
For all $\Lambda$’s, we observe a predominant regime of an exponential decay with $k$, shown with a dashed orange lines in Fig. 4(a,b). We anticipate a decay following the mixing argument of the preceding section. This exponential decay is typically cleaner for higher $\Lambda$’s (compare $k\in(10,45)$ in Fig. 4(a) with $k\in(70,250)$ in Fig. 4(b)), and lasts until we reach the numerical error cutoff for the KLD estimator, $\simeq 0.05$. Thus, for each $\Lambda$, we fit an exponential (as in Figs. 4(a,b)) to obtain a decay rate defined as $1/K_{\mathrm{m}}(\Lambda)$. We interpret these $K_{\mathrm{m}}$’s as the mixing-kick number. We verify for various $\Lambda$’s that it does not depend on the initial condition (by simulating $u=5\cdot 10^{-4}$ as well) or the sample size (by simulating $N=100$, too); $K_{\mathrm{m}}(\Lambda)$ remained identical within $3\%$ for all $\Lambda$’s, indicating that this convergence kick number is a robust quantity intrinsic to the dynamical system studied.
In Fig. 4(c) we plot the mixing-kick number $K_{\mathrm{m}}$ versus $\Lambda$. For the cubic map, we numerically find that $K_{\mathrm{m}}$ strongly depends on $\Lambda$, varying across two orders of magnitude in a manner consistent with a power law, $K_{\mathrm{m}}(\Lambda)\sim\Lambda^{-1.491\pm 0.046}$. This implies that the more violent the phase map, [^8] By more violent we mean a larger gain parameter $A$ of the cubic map, Eq. (2). Then, when $\Lambda>0$, $\Lambda(A)$ is a monotonously increasing function for the cubic map; see also [12, Fig. 15.3]. the faster the mixing in the considered dynamics. We caution that the exponential fit for the decay of the KLD with low $\Lambda$ is not perfect (specifically, for $\Lambda\lesssim 0.2$), so their $K_{\mathrm{m}}$ should be taken with a grain of salt.
We have no quantitative argument as to why $K_{\mathrm{m}}$ should diverge as $\Lambda\to 0^{+}$ as our findings suggest. Such a divergence is not obvious and should not apply generally, since the mixing property governing $K_{\mathrm{m}}$ is in principle different from the spreading property that governs $\Lambda$. We return to this point in Sec. V. In any case, the exact functional form of $K_{\mathrm{m}}(\Lambda)$ is not central to what follows; we will use the fact that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0^{+}$ as a qualitative guide.
To summarize, for $\Lambda>0$, the two phase statistics become identical after iterating sufficiently longer than the mixing-kick number $K_{\mathrm{m}}$. Thus, for the purpose of establishing synchronization, we will first make the reasonable assumption that both agents have received a number of common kicks exceeding an agreed threshold. In the next section we demonstrate this synchronization.
@@ IV Effective synchronization
In Sec. III, we found that after sufficiently many kicks, agents $\mathrm{A}$} and $\mathrm{B}$} effectively share $q^{(k)}(\varphi)$ in both the $\Lambda<0$ (synchronizing) and the $\Lambda>0$ (unsynchronizing) regimes. With $\Lambda<0$, the two agents agree at every moment (sufficiently-long after the most recent kick) on a specific phase to experimental accuracy and can use it to perform simultaneous actions or communicate. For $\Lambda>0$, while the resultant distributions do not asymptotically approach a sharp peak, the distributions of two agents become equivalent as numerically captured by the approach of the KLD among them to zero. The synchrony in the former and convergence in the latter remain with arbitrarily more kicks. Thus, any function of the sampled phases $\{\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}\}$ would be equal for both agents up to a statistical error. Below, we will consider a particular phase function we refer to as the “fiducial phase”, denoted $\varphi^{(k)}_{\mathrm{f}}=\varphi_{\mathrm{f}}[\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}$]. By phase function we mean that if $\{\varphi_{n}^{(k)}\}$ shift by $\Delta\varphi$, it must also shift by $\Delta\varphi$.
In order for agent $\mathrm{A}$} to act as though synchronized with agent $\mathrm{B}$}, it must be able to infer agent $\mathrm{B}$}’s fiducial phase $\varphi_{{\mathrm{f}},{\mathrm{B}}}$ at any given moment, given its phases $\{\varphi_{{\mathrm{A}}n}^{(k)}\}$. If agent $\mathrm{A}$}’s oscillator at that moment is $\varphi^{*}_{\mathrm{A}}$, its own fiducial phase $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}$ at that moment may be determined, since all phases advance at the same rate; see illustration in SI Appendix, Fig. S1. Specifically, $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{*}_{\mathrm{A}}={\rm const}=\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{(k)}_{\mathrm{A}}$; the same applies to agent $\mathrm{B}$}. Therefore, it remains to be shown that one may construct a function $\varphi^{(k)}_{{\mathrm{f}}}$ such that the estimated $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}=\varphi^{*}_{{\mathrm{f}},{\mathrm{B}}}$ to a good accuracy. We demonstrate below that this type of synchronization can indeed be achieved.
@@
The uncertainty in $\varphi_{\mathrm{f}}$ must depend on the degree of nonuniformity of $q(\varphi)$; surely, no $\varphi_{\mathrm{f}}$ can be unambiguously determined if $q(\varphi)$ is completely uniform. Further, uncertainty in measuring $q(\varphi)$ leads to further uncertainty in $\varphi_{\mathrm{f}}$. Below we test the discrepancy between $\varphi_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi_{{\mathrm{f}},{\mathrm{B}}}$ using the finite sampling method employed above. Our proof-of-concept example below shows that with a fixed sample size $N$, it is indeed feasible to attain small discrepancies which decrease with bigger $N$. These discrepancies depend strongly on the nonuniformity of the $q(\varphi)$ at hand, as measured by its entropy, Eq. (6). Thus, reliably small discrepancies require phase maps $\psi(\varphi)$ with small typical entropies.
@@
To determine a peak position, one needs a smooth estimate of $q(\phi)$ from the sampled values $\{\varphi_{1},\ldots,\varphi_{N}\}$. We obtain an estimate denoted $Q(\varphi,\sigma)$ smoothed to a “bandwidth” $\sigma$ using the method of kernel-density estimation [39, 40, 41], detailed in Appendix A.2. The reliability of the $\varphi_{\mathrm{f}}$ obtained depends on $\sigma$. Large $\sigma$ produces a $Q(\varphi;\sigma)$ with a few broad peaks whose positions have high uncertainty, whereas small $\sigma$ gives numerous statistically insignificant peaks; see illustration in SI Appendix, Figs. S7.999Finding an optimal bandwidth $\sigma$, particularly for highly-multimodal distributions, is an open question in the field [42, 43, 44, 45].
@@
This crude choice turns out to be sufficient for our purpose of demonstrating effective synchronization.